Cambridge A Level Mathematics 9709 — 2012 May/June Paper 6 · Variant 2

9709/62/M/J/12 · 6 questions · 50 marks · ≈56 min

The question paper and its mark scheme, free to read here and free to download. This is Cambridge’s own paper, exactly as it was sat.

← All Mathematics papersWhat was in this paper?

Question paper4 pages

Cambridge A Level Mathematics 9709 2012 May/June Paper 6 · Variant 2 question paper, page 1 of 4
Page 1 of 4
Cambridge A Level Mathematics 9709 2012 May/June Paper 6 · Variant 2 question paper, page 2 of 4
Page 2 of 4
Cambridge A Level Mathematics 9709 2012 May/June Paper 6 · Variant 2 question paper, page 3 of 4
Page 3 of 4
Cambridge A Level Mathematics 9709 2012 May/June Paper 6 · Variant 2 question paper, page 4 of 4
Page 4 of 4

Mark scheme6 pages

Answers below. Sit the paper first if you are practising.

Mark scheme, page 1 of 6
Page 1 of 6
Mark scheme, page 2 of 6
Page 2 of 6
Mark scheme, page 3 of 6
Page 3 of 6
Mark scheme, page 4 of 6
Page 4 of 6
Mark scheme, page 5 of 6
Page 5 of 6
Mark scheme, page 6 of 6
Page 6 of 6

Questions as text

Q1 · The ages, x years, of 150 cars are summarised by Σx = 645 and Σ x2 = 8287.5

1 The ages, x years, of 150 cars are summarised by Σx = 645 and Σ x2 = 8287.5. Find Σ(x −x)2, where x denotes the mean of x. [4]

Mark scheme: 1 x = 3.4 B1 4.3 or 645/150 or 18.49 seen M1 Subst in correct formula to find sd or var or  82875. 2  sd =  − 3.4  = 36.76 = 6.063 expand Σ ( x − x 2) correctly and substitute  150  Σ ( x − x 2) = 150× .6 063 2 M1 Mult by 150 = 5514 (5510) A1 [4] Answer rounding to 5510

More questions on Representation of data

Q2 · The random variable X has the probability distribution shown in the table

2 The random variable X has the probability distribution shown in the table. x 2 4 6 P(X = x) 0.5 0.4 0.1 Two independent values of X are chosen at random. The random variable Y takes the value 0 if the two values of X are the same. Otherwise the value of Y is the larger value of X minus the smaller value of X. (i) Draw up the probability distribution table for Y. [4] (ii) Find the expected value of Y. [1]

Mark scheme: 2 (i) B1 0, 2, 4 only seen for Y no probs needed. y 0 2 4 Accept other vals if P(value) = 0 seen in P(Y = y) 0.42 0.48 0.1 table, allow 0002244 with probs M1 Summing two or more 2-factor probs (can be implied) A1 One correct prob A1 [4] Correct table or list (ii) 0.96 + 0.4 = 1.36 B1ft [1] Ft their table for Y or X Σp = 1

More questions on Discrete random variables

Q3 · In Restaurant Bijoux 13% of customers rated the food as ‘poor’, 22% of customers rated…

3 In Restaurant Bijoux 13% of customers rated the food as ‘poor’, 22% of customers rated the food as ‘satisfactory’ and 65% rated it as ‘good’. A random sample of 12 customers who went for a meal at Restaurant Bijoux was taken. (i) Find the probability that more than 2 and fewer than 12 of them rated the food as ‘good’. [3] On a separate occasion, a random sample of n customers who went for a meal at the restaurant was taken. (ii) Find the smallest value of n for which the probability that at least 1 person will rate the food as ‘poor’ is greater than 0.95. [3]

Mark scheme: 3 (i) P(2 < X < 12) = 1 – P(0, 1, 2, 12) M1 Using binomial with 12Csomething and powers summing to 12, Σp = 1 = 1 – (0.35)12 – (0.65)(0.35)1112C1 – A1 Correct unsimplified answer (0.65)2(0.35)1012C2 – (0.65)12 = 1 – 0.0065359 = 0.993 A1 [3] Accept 0.994 from correct working only (ii) 1 – (0.87)n > 0.95 M1 Equality or inequality in (0.87 or 0.78 or 0.35), power n or n – 1, 0.95 or 0.05 0.05 > (0.87)n M1 Attempt to solve an equation with a power in (can be implied) n = 22 A1 [3] Correct answer

More questions on Probability

Q4 · The back-to-back stem-and-leaf diagram shows the values taken by two variables A and B

4 The back-to-back stem-and-leaf diagram shows the values taken by two variables A and B. A B (3) 3 1 0 15 1 3 3 5 (4) (2) 4 1 16 2 2 3 4 4 5 7 7 7 8 (10) (3) 8 3 3 17 0 1 3 3 3 4 6 6 7 9 9 (11) (12) 9 8 8 6 5 5 4 3 2 1 1 0 18 2 4 7 (3) (8) 9 9 8 8 6 5 4 2 19 1 5 (2) (5) 9 8 7 1 0 20 4 (1) Key: 4 16 7 means A = 0.164 and B = 0.167. (i) Find the median and the interquartile range for variable A. [3] (ii) You are given that, for variable B, the median is 0.171, the upper quartile is 0.179 and the lower quartile is 0.164. Draw box-and-whisker plots for A and B in a single diagram on graph paper. [3]

Mark scheme: 4 (i) A: median = 0.186, B1 IQ range = 0.198 – 0.179 M1 Subt LQ from their UQ = 0.019 A1ft [3] Correct IQ range ft dp in wrong place (ii) A B1ft 2 correct boxes ft (i) OK if superimposed B B1 2 pairs correct whiskers lines up to box not inside 0.15 0.16 0.17 0.18 0.19 0.20 0.21 B1 [3] Correct uniform scale from at least 0.15 to 0.21 seen. No scale no marks (ii) unless perfect A and B with all 10 values shown GCE AS/A LEVEL – May/June 2012 9709 62

More questions on Representation of data

Q5 · An English examination consists of 8 questions in Part A and 3 questions in Part B

5 An English examination consists of 8 questions in Part A and 3 questions in Part B. Candidates must choose 6 questions. The order in which questions are chosen does not matter. Find the number of ways in which the 6 questions can be chosen in each of the following cases. (i) There are no restrictions on which questions can be chosen. [1] (ii) Candidates must choose at least 4 questions from Part A. [3] (iii) Candidates must either choose both question 1 and question 2 in Part A, or choose neither of these questions. [3]

Mark scheme: 5 (i) 11C6 = 462 B1 OR A3 B3 or A4 B2 or A5 B1 or A6 = 8C3 + 8C4 × 3C2 + 8C5 × 3C1 + 8C6 B1 = 56 + 210 + 168 + 28 = 462 [1] (ii) 8C4 × 3C2 + 8C5 × 3C1 + 8C6 M1 ∑ 2 or more two-factor terms, P or C any numbers = 210 + 168 + 28 B1 Any correct option unsimplified = 406 A1 [3] Correct answer (iii) 9C4 + 9C6 = 126 + 84 M1 Summing 9Cx + 9Cy can be mult by 2 no other terms B1 126 or 84 seen or unsimplified 9C4, 9C6 = 210 A1 Correct answer OR 1,2 in A tog with : A1B3 + A2B2 + A3B1 + M1 ∑ 5 or more 2-factor 6Px or 6Cx with 3Cx or A4B0 + 1,2 out of A : A3B3 + A4B2 + 3Px only (can be mult by 2) A5B1 + A6B0 = 6C1 + 6C2 × 3C2 + 6C3 × 3C1 + 6C4 + 6C3 × B1 3 or more correct unsimplified options 3C3 + 6C4 × 3C2 + 6C5 × 3C1 + 6C6 = 6 + 45 + 60 + 15 + 20 + 45 + 18 + 1 = 210 A1 Correct answer OR 462 – 9C5 – 9C5 M1 subt two 9Cx options from their (i) B1 9C5 seen oe if using this method = 210 A1 [3] Correct answer

More questions on Permutations and combinations

Q6 · A box of biscuits contains 30 biscuits, some of which are wrapped in gold foil and some…

6 A box of biscuits contains 30 biscuits, some of which are wrapped in gold foil and some of which are unwrapped. Some of the biscuits are chocolate-covered. 12 biscuits are wrapped in gold foil, and of these biscuits, 7 are chocolate-covered. There are 17 chocolate-covered biscuits in total. (i) Copy and complete the table below to show the number of biscuits in each category. [2] Wrapped in gold foil Unwrapped Total Chocolate-covered Not chocolate-covered Total 30 A biscuit is selected at random from the box. (ii) Find the probability that the biscuit is wrapped in gold foil. [1] The biscuit is returned to the box. An unwrapped biscuit is then selected at random from the box. (iii) Find the probability that the biscuit is chocolate-covered. [1] The biscuit is returned to the box. A biscuit is then selected at random from the box. (iv) Find the probability that the biscuit is unwrapped, given that it is chocolate-covered. [1] The biscuit is returned to the box. Nasir then takes 4 biscuits without replacement from the box. (v) Find the probability that he takes exactly 2 wrapped biscuits. [4]

Mark scheme: 6 (i) wrapped unwrapped total choc 7 10 17 B1 One correct row or column numbers not choc 5 8 13 total 12 18 30 B1 [2] All correct including labels (ii) 12/30 (0.4) B1ft [1] Ft their table (iii) 10/18 (5/9) (0.556) B1ft [1] Ft their table (iv) 10/17 (0.588) B1ft [1] Ft their table (v) P(2 wrapped) = 12/30 × 11/29 × 18/28 × 17/27 × 4C2 M1 Mult by 4C2 M1 12 × 11 × 18 × 17 seen in num M1 30 × 29 × 28 × 27 seen in denom = 0.368 (374/1015) A1 Correct answer OR (12C2 × 18C2)/30C4 M1 12C2 seen mult or alone in num (not added) M1 18C2 seen mult or alone in num (not added) M1 30C4 seen in denom = 0.368 A1 [4] Correct answer GCE AS/A LEVEL – May/June 2012 9709 62  42 − 411.  M1 Standardising no cc no sq rt no sq 7 (i) P(> 42) = P  z >   4.3  = P(z > 0.2647) = 1 – 0.6045 = 0.3955 A1 Correct prob rounding to 0.395 or 0.396 Prob = (0.3955)(0.6045)23C1 M1 Binomial 3Cx powers summing to 3, any p, Σp = 1 = 0.433 or 0.434 A1 [4] Rounding to correct answer

More questions on Probability

What was in this paper

The subtopics covered by these 6 questions, and how many questions each got. Open one in a new tab to see every Cambridge question on it.

What you needed in this session

Cambridge’s own grade thresholds for 2012 May/June, Paper 6 · Variant 2. A higher threshold means an easier paper — the bar moves with how the cohort did.

A38/50
B32/50
E14/50